### v2

# Power law $\alpha$-Starobinsky inflation

Abstract
In this work we consider a generalization of Starobinsky inflation obtained by combining power law ($R^{\beta}$), and $\alpha$-Starobinsky inflation ($E$-model). The Einstein frame potential for this model is that of power law Starobinsky inflation modified by a parameter $\alpha$ in the exponential. After computing power spectra for scalar and tensor perturbations numerically, we perform MCMC analysis to put constraints on the potential parameter $\alpha$, $\beta$ and $M$, and the number of e-foldings $N_{pivot}$ during inflation, using Planck-2018, BICEP/Keck (BK18) and other LSS observations. We find $\log_{10}\alpha=0.37^{+0.82}_{-0.85}$, $\beta=1.969^{+0.020}_{-0.023}$, $M=\left(3.54^{+2.62}_{-1.73}\right)\times 10^{-5}$ and $N_{pivot}=47\pm{10}$. With these mean values of the potential parameters $\alpha$ and $\beta$, and varying $N_{pivot}$ between $40$ to $55$, we also find that the $r-n_{s}$ predictions of our model lie well within the $1\sigma$ bounds of joint constraints from combined analysis of ACT, Planck-2018, BICEP and BAO observations. We compute the Bayesian evidences for our proposed model, power law Starobinsky inflation, $\alpha$-Starobinsky inflation and Starobinsky inflation. Considering the Starobinsky model as the base model, we calculate the Bayes factor and find that our proposed model is preferred by the CMB and LSS observations.
I INTRODUCTION
Inflation Guth:1980zm refers to a period of rapid expansion in the early Universe before big-bang nucleosynthesis, proposed to address several challenges of the standard big bang cosmology such as the horizon problem, flatness problem, and the puzzling absence of magnetic monopoles Linde:1981mu ; Starobinsky:1980te . Inflation not only provides an explanation for these issues but also turns them into natural outcomes of the theory. It is achieved by a scalar field, named as inflaton, whose potential energy dominates the energy density of the universe causing quasiexponantial expansion of the universe for a very short period of time. Inflation produces density perturbations due to quantum fluctuations in the scalar field that are coupled to fluctuations in the metric, and tensor perturbations (primordial gravitational waves), which are generated due to the quantum fluctuations in the spacetime geometry. These quantum fluctuations serve as the seeds for the large scale structure of the Universe and cosmic microwave background anisotropy Mukhanov:1981xt ; Starobinsky:1982ee ; Guth:1985ya . The predictions of inflation i.e., adiabatic, nearly-scale invariant and Gaussian perturbations are firmly verified with the various CMB and LSS observations such as COBE COBE:1992syq , WMAP WMAP:2010qai , Planck Planck:2015sxf ; Planck:2018jri and BICEP-Keck array BICEP:2021xfz .
While inflation is generally seen as a successful idea, we still have not found a unique model of inflation that fits well within observational limits which have become more precise overtime. Since its invent, many different inflationary models have been proposed and examined Martin:2006rs ,Martin:2013tda with observational data. Analysis of recent Planck 2018 results Planck:2018jri shows that the simplest classic inflaton potentials, $\frac{1}{2}m^{2}\phi^{2}$ and $\lambda\phi^{4}$, are strongly ruled out. As proposed by Starobinsky Starobinsky:1980te , inflation can also be achieved without scalar field by adding $\frac{1}{M}R^{2}$ term, $R$ being the Ricci scalar, in the Einstein-Hilbert action. Named as Starobinsky inflation, this model is in remarkable agreement with Planck-2018 Planck:2018jri and BICEP/Keck BICEP:2021xfz ; Tristram:2021tvh observations as it predicts lower values for tensor-to-scalar ratio $r$. Originally described in the Jordan frame, Starobinsky inflation can be reinterpreted in the Einstein frame through a conformal transformation, where it looks like a single field inflationary model with a potential that flattens out at large field values ensuring a slow and sustained period of inflation. Interestingly this potential in the Einstein frame can be easily derived from no-scale supergravity with a non-compact $SU(2,1)/SU(2)\times U(1)$ symmetry Ellis:2013xoa , where we have a modulus field fixed by the other dynamics and the inflaton field is a part of the chiral superfield with a minimal Wess-Zumino superpotential. This enables incorporating inflation with various other aspects of particle physics phenomenology Ellis:2016ipm ; Ellis:2017jcp ; Ellis:2019opr . However, analysis of recently reported Atacama Cosmology Telescope data ACT:2025fju , in combination with Planck 2018, BAO and DESY Y1 data, indicates that the Starobinsky model is disfavored at the $2\sigma$ level ACT:2025tim .
Several generalizations of Starobinsky inflation have been proposed, such as, power law Starobinsky inflation having $R^{\beta}$ term in the Einstein-Hilbert action Schmidt:1989zz ; Maeda:1988ab ; Muller:1989rp ; Gottlober:1992rg ; DeFelice:2010aj ; Nojiri:2010wj ; Nojiri:2017ncd , which is based on higher-order metric theories of gravity, and $\alpha$-Starobinsky inflation ($E$-model) based on supergravity inspired deformation of the inflaton potential in the Einstein frame Ellis:2013nxa ; Ferrara:2013rsa ; Kallosh:2013yoa . Although inspired by higher-order metric theories of gravity, the potential for the power law Starobinsky inflation can also be derived from no-scale supergravity Chakravarty:2014yda . Power law Starobinsky inflation gained traction in 2014 when BICEP2 reported a large value of tensor-to-scalar ratio $r=0.2^{+0.07}_{-0.05}$ BICEP2:2014owc , as it was shown that this model generates a large $r$ compared to Starobinsky inflation for $\beta$ slightly less than $2$ Codello:2014sua ; Costa:2014lta ; Martin:2014lra . The primordial origin of BICEP2 signal of $B$-mode polarization was ruled out later and it was found the it is due to the foreground dust emission Planck:2014dmk . Observational constraints on $R^{\beta}$ inflation have been studied in Chakravarty:2014yda ; Motohashi:2014tra ; Odintsov:2022bpg ; Meza:2021xuq ; Saini:2023uqc and it is found that the current data allows a slight deviation from $\beta=2$ (see Saini:2023uqc for a detailed statistical analysis with Planck-2018 and BICEP3/Keck array data). In $\alpha$-Starobinsky inflation the scalaron potential in the Einstein frame contains a parameter $\alpha$ in the exponential that modifies the predictions for $n_{s}$ and $r$. The potential for this model can be obtained by generalizing the coefficient of the logarithm of the volume modulus field in the Kähler potential in $SU(2,1)/SU(2)\times U(1)$ no-scale supergravity with a suitable choice of the superpotential having both the volume modulus field and the chiral superfield. The parameter $\alpha$ is these models is related to the geometry of the Kähler manifold. The observational constraints on $\alpha$-Starobinsky inflation from Planck and BICEP/Keck array data are studied in Planck:2015sxf ; Planck:2018jri ; Ueno:2016dim ; Ellis:2020xmk ; Ellis:2021kad ; ElBourakadi:2022lqf ; Sarkar:2021ird ; Saini:2024mun , and it is found that the current data is not sufficient to constrain the parameter $\alpha$. However, an upper limit on $\alpha$ i.e., $\alpha\leq 39.81$ with $95\%$ C. L. is obtained in Saini:2024mun from Planck-2018 and BICEP/Keck array observations by performing a detailed statistical analysis. A two field inflation model with both the field having $\alpha$-attractor potentials have also been studied in Rodrigues:2020fle . The $E$-model potential in brane inflation has also been studied in Sabir:2019wel . $R^{2}$ attractors and $\alpha$-attractors inflationary models as a subclass of $f(R)$ gravity in the Jordan frame have been studied in Odintsov:2020thl (see also Odintsov:2023ypt for various phenomenological aspects of these models). In SantosdaCosta:2020dyl the variations from Starobinsky potential in the Einstein frame have been studied based on a potential derived from brane inflation. Extensions of Starobinsky model with $R^{3}$, $R^{4}$ and $R^{3/2}$, in addition to $R^{2}$ term, have also been studied in Ivanov:2021chn in the light of Planck 2018 observations. It is shown in Gialamas:2025ofz that the Starobinsky inflation with an additional $R^{3}$ term is in full agreement with recent ACT observations.
In this work we investigate a potential obtained by the combination of power law and $\alpha$-Starobinsky model in the Einstein frame having two parameters $\alpha$ and $\beta$ to analyze the deviation from the Starobinsky inflation. Since the required conformal transformations from Jordan frame to Einstein frame are related to the Kähler potential Cremmer:1978hn ; Cremmer:1982en , we consider $R^{\beta}$ term in the Einstein-Hilbert action and derive the Einstein frame potential by introducing the parameter $\alpha$ in the conformal transformations. We also provide a derivation of the potential in the no-scale supergravity framework as done in Chakravarty:2014yda with the Kähler potential having the parameter $\alpha$ as the coefficient of the logarithm of the volume modulus field. We employ ModeChord 2021ascl an updated version of ModeCode Mortonson:2010er to solve the background and perturbation equations for the inflaton without assuming the slow-roll approximation, enabling the computation of primordial power spectra for scalar and tensor perturbations. These power spectra are used in CAMB Lewis:1999bs to generate angular power spectra for CMB anisotropies and polarization. We then perform MCMC analysis using CosmoMC Lewis:2002ah to directly constrain the parameters of inflaton potential and number of e-foldings using Planck-2018, BICEP/Keck array and other LSS observations.
We also calculate Bayesian evidence for power law model, $\alpha$-Starobinsky model, power law $\alpha$-Starobinsky model and Starobinsky model using the MCEvidence Heavens:2017afc . which is a tool in Python that computes the Bayesian evidences using data from MCMC simulations. It uses a method called $k$-th nearest neighbor to estimate how likely each model is, which is important when comparing different models in Bayesian analysis. By computing the Bayes factors it can be determined that how well these generalized models align with or outperform the original Starobinsky model in the light of current CMB and LSS observations.
The rest of this paper is organized as follows. In section II, we begin with brief discussion of the Power law $\alpha$ Starobinsky model. Section III provides a comprehensive overview of the background and perturbation equations used in ModeChord to compute the primordial power spectra. We present the results of testing the inflationary models against observational data using Markov chain Monte Carlo (MCMC ) analysis in section IV and in section V, we describe the insights gained through Bayesian evidence calculations. Finally, in Section VI, we summarize our conclusions.
II Power law $\alpha$-Starobinsky Model
The modified Starobinsky model, which includes an $R^{\beta}$ correction, has been analyzed in DeFelice:2010aj ; Chakravarty:2014yda ; Motohashi:2014tra . The corresponding action in Jordan frame is given by,
$S_{J}=\frac{-M_{Pl}^{2}}{2}\int\sqrt{-g}f(R)d^{4}x,\quad\quad f(R)=\left(R+\frac{1}{6M^{2}}\frac{R^{\beta}}{M_{Pl}^{2\beta-2}}\right),$ (1)
where $M_{Pl}^{2}=(8\pi G)^{-1}$, g is the determinant of the metric $g_{\mu\nu}$, and $M$ is a dimensionless real parameter. We will take $M_{Pl}=1$ in the the rest of the paper.
To obtain the action in the Einstein frame we use conformal transformations $\tilde{g}_{\mu\nu}(x)=\Omega(x)g_{\mu\nu}(x)$, where a tilde denotes the quantities in the Einstein frame. The Ricci scalar $R$ defined in the Jordan frame can be expressed in terms of the Einstein frame Ricci scalar $\tilde{R}$ via the conformal transformation connecting the two frames as
$R=\Omega\left(\tilde{R}+3\widetilde{\Box}\omega-\frac{3}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\omega}\partial_{\nu}{\omega}\right),$ (2)
here $\omega\equiv\ln\Omega$, $\widetilde{\Box}\omega\equiv\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\left(\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}{\omega}\right)$ and $\partial_{\mu}{\omega}=\frac{\partial\omega}{\partial\tilde{x}^{\mu}}$.
To obtain action in the Einstein frame we choose $\Omega=F=\frac{\partial f(R)}{\partial R}$ and also introduce a new scalar field $\chi$ defined as
$\chi\equiv\frac{\sqrt{6\alpha}}{2}\ln F.$ (3)
This leads to
$\Omega=\exp\left(\frac{2\chi}{\sqrt{6\alpha}}\right).$ (4)
Here, the parameter $\alpha$ is a dimensionless parameter, and it is introduced to obtain the potential of $\alpha$-Starobinsky inflation in the Einstein frame. Now, using Eq. (2) and (4) along with $\sqrt{-g}=\Omega^{-2}\sqrt{-\tilde{g}}$, the action (1) can be transformed to the Einstein-Hilbert form as
$S_{E}=\int d^{4}x\sqrt{-\tilde{g}}\left(-\frac{1}{2}\tilde{R}+\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\chi}\partial_{\nu}{\chi}+V(\chi)\right),$ (5)
where the Einstein frame potential $V(\chi)$ is given as
$V(\chi)=\frac{(RF(R)-f(R))}{2F(R)^{2}}.$ (6)
Thus, the potential for the power law $\alpha$-Starobinsky Model can be expressed as
$V({\chi})=(\frac{{\beta}-1}{2})\left(\frac{6M^{2}}{{\beta}^{\beta}}\right)^{\frac{1}{\beta-1}}\exp\left[{\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{2-\beta}{\beta-1}\right)}\right]\times\left(1-\exp\left(-\frac{2\chi}{\sqrt{6\alpha}}\right)\right)^{\frac{\beta}{\beta-1}}.\\$ (7)
We can see that this potential reduces to the Starobinsky ($R^{2}$) potential in the Einstein frame for $\beta=2$ and $\alpha=1$. In the large field limit $\chi\gg 2\sqrt{6}$ and $1<\beta<2$ the potential from Eq. (7) simplifies to
$V({\chi})=(\frac{{\beta}-1}{2})\left(\frac{6M^{2}}{{\beta}^{\beta}}\right)^{\frac{1}{\beta-1}}\exp\left[{\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{2-\beta}{\beta-1}\right)}\right].$ (8)
As there is a correspondence between the $R^{2}$ (de Sitter) and $R+R^{2}$ (Starobinsky) theories of gravity and no-scale supergravity Ellis:2017xwz , the supergravity Lagrangian written in the form of Einstein-Hilbert action requires a conformal transform expressed in terms of Kähler potential $\Omega^{2}=\exp(-K/3)$ Cremmer:1978hn ; Cremmer:1982en . As the potential for $\alpha$-Starbinsky inflation can be obtained by changing the Kähler curvature $\alpha\neq 1$, the corresponding conformal transformation can be expressed in terms of Kähler potential as $\Omega^{2}=\exp(-K/(3\alpha))$ Ellis:2019bmm . This justifies the introduction of parameter $\alpha$ in Eq. (3). A detailed derivation of the potential (8) from no-scale supergravity is provided in the next section.
II.1 No-scale supergravity realization of the power law $\alpha$-Starobinsky Model
The Einstein frame potential for the generalized power law $\frac{1}{M^{2}}R^{\beta}$ model has been obtained in Chakravarty:2014yda by incorporating a $(\Phi+\Phi^{*})^{n}$ term into the minimal no-scale Kähler potential along with a Wess-Zumino superpotential Ellis:2013xoa ; Chakravarty:2014yda ; Ellis:2020xmk . The potential for power law $\alpha$-Starobinsky model can be obtained by generalizing the coefficient of the volume modulus field in the Kähler potential and following the same approach as in Chakravarty:2014yda . So the Kähler potential with parameter $\alpha$ is given as
$K=-3\alpha\ln\left(T+T^{*}-\frac{1}{2}(\phi+\phi^{*})^{n}\right),$ (9)
here, the field $T$ represents a modulus field and $\phi$ is a matter field that serves as the inflaton. The parameter $\alpha$ is inversely related to the curvature of the Kähler manifold ($R=\frac{2}{3\alpha}$), and is known as Kähler curvature parameter. A vacuum expectation value (vev) is assigned to the $T$ field such that $\langle T+T^{*}\rangle=2\langle\text{Re}\,T\rangle=c>0,\quad\langle\text{Im}\,T\rangle=0$ and the imaginary component of the inflaton field $\phi$ is set to zero. $\phi=\phi^{*}$. So the effective Lagrangian in the Einstein frame becomes
$\mathcal{L}_{E}=\frac{n\alpha(2\phi)^{n-2}\left[c(n-1)+\frac{(2\phi)^{n}}{12}\right]}{4\left[c-\frac{(2\phi)^{n}}{12}\right]^{2}}(\partial_{\mu}\phi)^{2}-\frac{4(2\phi)^{2-n}}{\alpha n(n-1)\left[c-\frac{(2\phi)^{n}}{12}\right]^{3\alpha-1}}\left|\frac{\partial W}{\partial\phi}\right|^{2}.$ (10)
To make the kinetic term canonical in the above Lagrangian, the field $\phi$ can be redefined to $\chi$ as
$\frac{\partial\chi}{\partial\phi}=\frac{\sqrt{n\alpha(2\phi)^{n-2}\left[c(n-1)+\frac{(2\phi)^{n}}{12}\right]}}{2\left(c-\frac{(2\phi)^{n}}{12}\right)}.$ (11)
By integrating above equation, we get
$\phi\approx\frac{1}{2}\exp\left(\frac{2\chi}{\sqrt{3n\alpha}}\right)\left[1+\frac{6c(n+1)}{n}\exp\left(-\frac{2n\chi}{\sqrt{3n\alpha}}\right)\right].$ (12)
Substituting above equation and Wess-Zumino superpotential $W=\frac{\mu}{2}\phi^{2}-\frac{\lambda}{3}\phi^{3}$ Ellis:2013xoa ; Chakravarty:2014yda ; Ellis:2020xmk in the potential term of Lagrangian $\mathcal{L}_{E}$, we get
$V(\chi)$ $=\frac{(12)^{3\alpha-1}\mu^{2}}{\alpha\,n(n-1)(-1)^{3\alpha-1}}\exp\left[\frac{2\chi}{\sqrt{6\alpha}}\cdot\frac{3\sqrt{2}(2-n\alpha)}{\sqrt{n}}\right]$
$\quad\times\left[1-\frac{2\mu}{\lambda}\exp\left(-\frac{2\chi}{\sqrt{3n\alpha}}\right)-\frac{9c(n^{2}-n-2)}{n}\exp\left(-\frac{2n\chi}{\sqrt{3n\alpha}}\right)\right]^{2}.$ (13)
We have assumed $1<n<2$, and in the large field limit $\chi\gg\sqrt{3n}$ above equation is equivalent to
$V=\frac{(12)^{3\alpha-1}\mu^{2}}{\alpha n(n-1)(-1)^{3\alpha-1}}\exp\left[\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{3\sqrt{2}(2-n\alpha)}{\sqrt{n}}\right)\right].$ (14)
By comparing the power law potential 8 and SUGRA potential 14, we can show the relation between the parameters of the two models.
Figure 1: Variation of potential (7) with the field for different values of $\alpha$ and $\beta$, as indicated. Both the potential and the field are expressed in $M_{Pl}=1$ units.
The plot 1 depicts the variation of potential (7) for different combinations of the parameters $\alpha$ and $\beta$. For $\beta=2$ and $\alpha=1$, the behavior is representative of the Starobinsky $R^{2}$ inflationary model. When $\beta$ deviates from 2:
[…]
The parameter $\alpha$ further modulates the steepness:
[…]
Overall, the interplay between $\alpha$ and $\beta$ governs the inflationary dynamics. This plot is crucial for exploring how these parameters influence observable predictions.
III Inflationary dynamics and power spectra
In this section we describe the background equations for inflationary dynamics and the perturbation equations, which are solved numerically using ModeCode. Here, the cosmic time derivatives will be indicated using an overdot notation, while differentiation with respect to number of e-foldings ($N$) will be denoted by a prime symbol. We will express all equations in natural units i.e. $c=\hbar=8\pi G=1.$
A. Background equations
To describe the inflationary dynamics in the Einstein frame, the Einstein-Hilbert action is given as
$S_{E}=\int d^{4}x\,\sqrt{|g|}\left(\frac{1}{2}{\tilde{R}}+\frac{1}{2}\nabla_{\mu}\chi\nabla^{\mu}\chi-V(\chi)\right),$ (15)
here $R$ and $\chi$ are the Ricci scalar and scalar field respectively, and $V(\chi)$ represents the potential of the scalar field.
The line element is given by
$ds^{2}=dt^{2}-a^{2}\delta_{ij}dx^{i}dx^{j},$
where $a$ is the scale factor. During inflation the expansion of the Universe is governed by the Friedmann equations
$\quad\quad\quad\quad H^{2}=\frac{1}{3}\left(\frac{1}{2}\dot{\chi}^{2}+V(\chi)\right),$ (16)
and
$\dot{H}=-\frac{1}{2}\dot{\chi}^{2},$ (17)
where $H$ is the Hubble parameter, $H=\frac{\dot{a}}{a}$.
The equation of motion for the $\chi$ field in the expanding universe is
$\ddot{\chi}+3H\dot{\chi}+\frac{dV}{d\chi}=0.$ (18)
Once we apply the initial conditions, the evolution of universe can be fully specified by any two of Eq. (16) to Eq. (18). These equations are collectively called as background equations. Since we use the number of e-foldings, $N=\ln a$, as an independent variable for numerical solutions, we can express the background equations in terms of $N$ as follows
$H^{2}$ $=$ $\frac{\frac{1}{3}V(\chi)}{1-\frac{1}{6}\chi^{\prime 2}},$ (19)
$H^{\prime}$ $=$ $-\frac{1}{2}H\chi^{\prime 2},$ (20)
and
$\chi^{\prime\prime}+\left(\frac{H^{\prime}}{H}+3\right)\chi^{\prime}+\frac{1}{H^{2}}\frac{dV(\chi)}{d\chi}=0.$ (21)
These Background equations can be solved numerically along with perturbation equations.
B. Mukhanov-Sasaki Equations
CMB observations indicate that there are small deviations from the ideal state of universe. The inhomogeneities observed in the CMB are of the order of $10^{-5}$, it is adequate for constraining analysis to linear perturbations of scalar field and metric. The scalar field $\chi$ is perturbed as follows:
$\chi(t,x)=\bar{\chi}(t)+\delta\chi(t,x),$ (22)
and the line element is given by
$ds^{2}=(1+2\chi)dt^{2}+2a(\partial_{i}B-S_{i})dx^{i}dt-a^{2}(1-2\Psi)\delta_{ij}dx^{i}dx^{j}-a^{2}\left(2\partial_{i}\partial_{j}E+2\partial_{(i}F_{j)}+h_{ij}\right)dx^{i}dx^{j}.$ (23)
In above equation, the perturbations have been decomposed into scalar, vector and tensor components. The SVT decomposition allows us to classify perturbations based on their transformation properties. The vector component is exponentially diluted in an exponentially expanding universe Mukhanov:1990me so the vector perturbations become negligible. The comoving curvature perturbation, $\mathcal{R}$, is the gauge-invariant quantity used to calculate the primordial power spectrum. It describes curvature of spatial hypersurfaces where $\chi$ remains invariant, given by:
$\mathcal{R}\approx\Psi+\frac{H}{\dot{\bar{\chi}}}\delta\chi.$ (24)
A general action for $\mathcal{R}$ is obtained by substituting Eq. (22) to Eq. (24) into Eq. (15) and we can write the second-order action for scalar perturbations during inflation in terms of the Mukhanov variable $u=z\mathcal{R}$ as Mukhanov:1988jd ; Sasaki:1986hm :
$S_{(2)}=\frac{1}{2}\int\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\left(\frac{du}{d\tau}\right)^{2}-\left(\partial_{i}u\right)^{2}+u^{2}\frac{1}{z}\frac{d^{2}z}{d\tau^{2}}\right].$ (25)
The Fourier expansion of variable $v$ expresses as:
$u(\tau,\mathbf{x})=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}\,u_{\mathbf{k}}(\tau)e^{i\mathbf{k}\cdot\mathbf{x}}.$
We obtain the equation of motion (Mukhanov-Sasaki equation) for the Mukhanov variable by varying the action described above.
$\frac{d^{2}u_{k}}{d\tau^{2}}+\left(k^{2}-\frac{1}{z}\frac{d^{2}z}{d\tau^{2}}\right)u_{k}=0.$ (26)
We can solve this second-order differential equation exactly only when we fully define the background dynamics, allowing us to know the evolution of $z$. Here, $z=\frac{1}{H}\frac{d\chi}{d\tau}$.
Mukhanov-Sasaki equation in terms of the parameter N can be rewritten as:
$u_{k}^{\prime\prime}+\left(\frac{H^{\prime}}{H}+1\right)u_{k}^{\prime}+\Biggl\{\frac{k^{2}}{a^{2}H^{2}}-\left[2-4\frac{H^{\prime}}{H}\frac{\chi^{\prime\prime}}{\chi^{\prime}}-2\left(\frac{H^{\prime}}{H}\right)^{2}-5\frac{H^{\prime}}{H}-\frac{1}{H^{2}}\frac{d^{2}V}{d\chi^{2}}\right]\Biggr\}u_{k}=0,$ (27)
$\frac{d^{2}v_{k}}{d\tau^{2}}+\left(k^{2}-\frac{1}{a}\frac{d^{2}a}{d\tau^{2}}\right)v_{k}=0,$ (28)
and
$v_{k}^{\prime\prime}+\left(\frac{H^{\prime}}{H}+1\right)v_{k}^{\prime}+\left[\frac{k^{2}}{a^{2}H^{2}}-\left(\frac{H^{\prime}}{H}+2\right)\right]v_{k}=0.$ (29)
C. Power spectra
We define the primordial power spectrum of curvature perturbations, P(k), as a function of the vacuum expectation value of R.
$\mathcal{P_{\mathcal{R}}}=\frac{k^{3}}{2\pi^{2}}\langle\mathcal{R}_{k}\mathcal{R}_{k^{\prime}}^{*}\rangle\delta^{3}(k-k^{\prime}),$ (30)
in this context, $\delta$ denotes the Dirac delta function, and the factor $\frac{2\pi^{2}}{k^{3}}$ is selected to obey the standard Fourier conventions. Consequently, $\mathcal{P_{\mathcal{R}}}(k)$ is connected to $u_{k}$ and $z$ through the following relation:
$\mathcal{P_{\mathcal{R}}}(k)=\frac{k^{3}}{2\pi^{2}}\bigg|\frac{u_{k}}{z}\bigg|^{2}.$ (31)
Similarly, the primordial tensor power spectrum is connected to $v_{k}$ and $z$ via
$\mathcal{P}_{t}(k)=\frac{4k^{3}}{\pi^{2}}\bigg|\frac{v_{k}}{a}\bigg|^{2}.$ (32)
The scalar spectral index $n_{s}$ and the tensor spectral index $n_{t}$ are extracted from the numerically obtained power spectra by using their definition as Bassett:2005xm .
$\quad\quad n_{s}=1+\frac{d\ln\mathcal{P_{\mathcal{R}}}}{d\ln k},$ (33)
and
$n_{t}=\frac{d\ln\mathcal{P}_{t}}{d\ln k}.$ (34)
Similarly, the tensor-to-scalar ratio $r$, defined by Bassett:2005xm
$r=\frac{\mathcal{P}_{t}}{\mathcal{P_{\mathcal{R}}}}.$ (35)
In our analysis both $n_{s}$ and $r$ are derived parameters, and the parameters of the potential (7) $M$, $\alpha$ and $\beta$ are directly constrained from CMB observations.
IV Observational constraints
[…]
Parameter Prior range
$\mathbf{N_{\text{pivot}}}$ [25 : 90]
$\mathbf{\log_{10}M}$ [-8.0 : -3.0]
$\mathbf{\beta}$ [1.90 : 2.07]
$\mathbf{\log_{10}\alpha}$ [-8.0 : 4.0]
Table 1: Priors on $N_{\text{pivot}}$ and model parameters for power law $\alpha$-Starobinsky model.
Table 2 shows the observational constraints on the parameters of power law $\alpha$-Starobinsky potential (7) and $N_{pivot}$, obtained from the Planck 2018 observations in combination with BAO and BICEP (BK18) BICEP:2021xfz .
Parameter 68% limits 95% limits 99% limits
$N_{\text{pivot}}$ $47^{+9}_{-4}$ $47\pm{10}$ $47^{+10}_{-20}$
$\log_{10}M$ $45^{+0.16}_{-0.12} 4.4$$5^{+0.24}_{-0.29} 4.4$$45^{+0.32}_{-0.34}$
$\beta$ $1.969^{+0.018}_{-0.0086}$ $1.969^{+0.020}_{-0.023}$ $1.969^{+0.022}_{-0.027}$
$\log_{10}\alpha$ $0.37\pm{0.43}$ $0.37^{+0.82}_{-0.85}$ $0.4^{+1.0}_{-1.0}$
$n_{s}$ $0.9685\pm{0.0031}$ $0.9685^{+0.0064}_{-0.0060}$ $0.9685^{+0.0084}_{-0.0081}$
[…]
Table 2: Constraints on parameters of potential, $r$ and $n_{s}$ using Planck-2018, BICEP/Keck (BK18) and BAO observations.
It can be seen from the table that the constraints on the parameters $\beta$ and $\alpha$ are
$\beta=1.969^{+0.020}_{-0.023},\,\,\,95\%\,C.\,L.,$ (36)
$\log_{10}\alpha=0.37^{+0.82}_{-0.85},\,\,\,95\%\,C.\,L.$ (37)
This indicates that the current observations allow the deviation from $\beta=2$ by $2\sigma$, while $\alpha=1$ is still consistent the observations. The number of e-foldings obtained for this model is
$N_{pivot}=47\pm 10\,\,\,95\%\text{C. L.}$ (38)
[…]
Fig. 2 illustrates the posterior probability distribution and the confidence regions at $68\%$ and $95\%$ confidence levels for the potential parameter and the derived parameters $(n_{s},r)$. From the Fig. 2, it can be seen that the parameters $(N_{pivot},\,\log_{10}\alpha)$, $(\log_{10}M,\,\log_{10}\alpha)$ and $(\log_{10}M,\,\beta)$ are strongly correlated, while $(\log_{10}M,\,N_{pivot})$, $(\beta,N_{pivot})$ are weekly correlated or almost uncorrelated. The contours of $(\beta,\log_{10}M)$ are elongated and tilted downward, indicating a negative correlation. As $\log_{10}M$ increases, $\beta$ decreases slightly. For the parameters $(\log_{10}\alpha$, $\log_{10}M)$ contours are stretched and tilted upward, showing a positive correlation. However, in case of power law Starobinsky inflation $(\beta,\,N_{pivot})$ and $(\log_{10}M,N_{pivot})$ are strongly correlated Saini:2023uqc , and in case of $\alpha$-Starobinsky inflation $N_{pivot}$ and $\alpha$ are not correlated Saini:2024mun . Fig. 2 shows the joint $68\%$ and $95\%$ C.L. constraints on scalar spectral index $n_{s}$ and tensor-to-scalar ratio $r$. These are derived parameters, and their constraints are determined by using Eqns. (33) and (35). As depicted in Fig. 2, the derived parameters $n_{s}$ and $r$ are almost uncorrelated, contrary to standard inflationary scenario and $\alpha$-Starobinsky inflation Saini:2024mun , but, similar to power law Starobinsky inflation Saini:2023uqc .
Fig. 3 shows the $r-n_{s}$ predictions for our model with mean values of the potential parameters along with joint $68\%$ and $95\%$ C.L. constraints from ACT, Planck, BAO from DESI and BICEP observations ACT:2025fju ; ACT:2025tim . We choose the parameters $\log_{10}\alpha=0.37$, $\beta=1.969$, $M=10^{-5}$, and vary number of e-foldings $N_{pivot}$ between $40$ to $55$ to determine the dependence of tensor-to-scalar ratio $r$ on $n_{s}$ shown by black line in the figure. It is clear from Fig. 3 that the $r-n_{s}$ predictions of our model agree within $1\sigma$ with the recent ACT observations.
Figure 3: The $r-n_{s}$ predictions for power law $\alpha$-Starobinsky inflation (black line) along with the joint constraints from ACT, Planck, BICEP/Keck and BAO observations ACT:2025tim .
V BAYESIAN MODEL SELECTION
[…]
$p(\theta|y,{\cal M})=\frac{\mathcal{L}(y|\theta,{\cal M})\pi(\theta|{\cal M})}{p(y|{\cal M})}.$ (39)
From above equation, the evidence is obtained by integrating the unnormalized posterior distribution over the parameter space.
$p(y|{\cal M})=\int d^{n}\theta\,\mathcal{L}(y|\theta,{\cal M})\pi(\theta|{\cal M}).$ (40)
[…]
$p({\cal M}|y)=\frac{p(y|{\cal M})\pi({\cal M})}{p(y)},$ (41)
[…]
$p(y)=\sum_{i}p(y|{\cal M}_{i})\pi({\cal M}_{i}).$ (42)
Typically, the comparison between two or more models is made by calculating the ratio of their evidences. The Bayes’ ratio
$B_{ij}=\ln\left(\frac{p(y|{\cal M}_{i})}{p(y|{\cal M}_{j})}\right),$ (43)
is subsequently interpreted according to Jeffreys’ guidelines Jeffreys:1939xee . The observational support for the underlying model ${\cal M}_{i}$ relative to ${\cal M}_{j}$ is quantified based on the values of $B_{ij}$(or alternatively $\ln B_{ij}$). Additionally, according to the ”revised Jeffreys’ scale” by Kass and Raftery Kass:1995loi , the values of $|\ln B_{ij}|$ in the intervals $(0,1.0)$, $(1.0,3.0)$, $(3.0,5.0)$, and $(5.0,\infty)$ indicate weak, definite/positive, strong and Very strong evidence, respectively.
Here we use the publicly available code MCEvidence Heavens:2017afc to calculate the Bayesian evidences for the models. MCEvidence uses only the MCMC chains produced by CosmoMC and provides the logarithm of the Bayes factor, $\ln B_{ij}$. This value is then used to evaluate whether model $M_{i}$ is preferred over model $M_{j}$.
Model $\ln B$ $\Delta\ln B_{ij}$
Starobinsky -1958.506 -
Power law -1957.047 1.459
$\alpha$-Starobinsky -1957.054 1.452
Power law $\alpha$-Starobinsky -1956.820 1.686
Table 3: The column $\Delta\ln B_{ij}$ corresponds to the comparison of Model Power law, Model $\alpha$-Starobinsky and Power law $\alpha$-Starobinsky with Starobinsky Model.
The Bayesian evidences and Bayes factors for various models are shown in Table. 3. We have computed Bayes factor by treating Starobinsky inflation as our reference. It can be seen from the Table that the power law $\alpha$-Starobinsky inflation has a higher Bayes factor compared to other models, hence it is preferred by current observations.
VI Conclusions
One of the best suited model from Planck 2018 Planck:2018jri and BICEP/Keck (Bk18) observations BICEP:2021xfz is Starobinsky inflation Starobinsky:1980te , where the inflation can be achieved without scalar field by adding $R^{2}$ term in the Einstein-Hilbert action. An interesting feature of Starobinsky inflation is that its potential in the Einstein frame can be realized in the framework of no-scale supergravity with a non-compact $SU(2,1)/SU(2)\times U(1)$ symmetry Ellis:2013xoa , where we have a modulus field fixed by the other dynamics and the inflaton field is a part of the chiral superfield with a minimal Wess-Zumino superpotential. There are generalizations of Starobinsky inflation such as power law, $R^{\beta}$, Starobinsky inflation based on higher order metric theories of gravity Schmidt:1989zz ; Maeda:1988ab ; Muller:1989rp ; Gottlober:1992rg ; DeFelice:2010aj ; Nojiri:2010wj ; Nojiri:2017ncd , and $\alpha$-Starobinsky inflation ($E$-model) based on supergravity Ellis:2013nxa ; Ferrara:2013rsa ; Kallosh:2013yoa , where the inflaton potential has a parameter $\alpha$ in the exponential in the Einstein frame. The Einstein frame potential for the power law Starobinsky inflation can also be derived from no-scale supergravity Chakravarty:2014yda .
In this work we examine a potential formulated by combining the power law and $\alpha$-Starobinsky model, Eq. (7), in the Einstein frame, which contains both the parameters $\alpha$ and $\beta$ to analyze the deviation from Starobinsky inflation in the light of current observation. We obtain the potential for power law $\alpha$-Starobinsky inflation by considering $R^{\beta}$ term in the Einstein action and introducing a parameter $\alpha$ in the conformal transformation via field redefinition motivated by supergravity Cremmer:1978hn ; Cremmer:1982en . We also derive the potential from no-scale supergravity with an underlying noncompact $SU(2,1)/SU(2)\times U(1)$. We compute the primordial power spectra numerically using ModeChord 2021ascl by solving background and perturbation equations without using the slow-roll approximation in the Einstein frame. With the help of this we perform MCMC analysis using CosmoMC to put constraints on the parameters $M$, $\alpha$ and $\beta$ of the potential (7) and the number of e-foldings $N_{pivot}$ from the end of inflation to the time when the scales corresponding to $k_{pivot}$ left the Hubble radius during inflation. We find $\log_{10}\alpha=0.37^{+0.82}_{-0.85}$ and $\beta=1.969^{+0.020}_{-0.023}$ at $95\%$ C.L. This indicates that the current CMB and LSS observations allow $2\sigma$ deviation from $\beta=2$. However, the limits on $\log_{10}\alpha$ are larger than the central value indicating that $\alpha=1$ is consistent with the current observations. These results are similar to power law Starobinsky inflation Saini:2023uqc and $\alpha$-Starobinsky inflation Saini:2024mun . We also find the parameter that determines the energy scale of inflation $M=\left(M=3.34^{+2.62}_{-1.73}\right)\times 10^{-5}$ at $95\%$ C. L, and the number of e-folds from the end of inflation to the time when the pivot scale $k_{pivot}$ leaves the inflationary horizon $N_{pivot}=47\pm 10$ at $95\%$ C. L. Interestingly, by performing MCMC analysis, we find that the parameters $\beta$ and $N_{pivot}$ are uncorrelated, which is in contrast to power law Starobinsky inflation Saini:2023uqc where these parameters are strongly correlated. Similarly, the parameters $\log_{10}\alpha$ and $N_{pivot}$ are strongly correlated in contrast to $\alpha$-Starobinsky inflation Saini:2024mun where these are uncorrelated. The parameters $\log_{10}M$ and $N_{pivot}$ are weakly correlated similar to $\alpha$-Starobinsky inflation Saini:2024mun , but, contrary to power law Starobinsky inflation Saini:2023uqc . There is also no correlation between the derived parameters $n_{s}$ and $r$ contrary to $\alpha$-Starobinsky inflation Saini:2024mun , but, similar to power law Starobinsky inflation Saini:2023uqc .
Recently the Atacama Cosmology Telescope, in combination with Planck 2018, BAO data from DESI Y1 ACT:2025fju , has reported $n_{s}=0.9743\pm 0.0034$, which is slightly higher than the value reported by Planck 2018 observations, i.e., $n_{s}=0.9651\pm 0.004$ Planck:2018jri . This suggests that the Starobinsky model is now disfavored at the $2\sigma$ level ACT:2025tim as it predicts lower values for spectral index. We also obtain $r-n_{s}$ predictions for our model for the mean values of the potential parameters $\alpha$ and $\beta$ and varying $N_{pivot}$ between $40$ to $55$. We find that these predictions lie with $68\%$ C.L. of joint constraints from ACT, Planck-2018, BICEP and BAO data from DESI Y1 ACT:2025fju ; ACT:2025tim .
To determine which of $R^{\beta}$, $\alpha$-Starobinsky, or power law $\alpha$-Starobinsky model is best suited by the observation, we calculate the Bayesian evidences of each model using MCEvidence Heavens:2017afc . We also calculate the logarithm of the Bayes factor, $\ln B_{ij}$ given in Eq. (40) for each of the three models. Here, $i$ represents $R^{\beta}$, $\alpha$-Starobinsky, and power law $\alpha$-Starobinsky model, while $j$ represents the Starobinsky model as reference model. So we can conclude from table 3 that power law $\alpha$-Starobinsky model has higher Bayesian evidence as a result it shows greater deviation from Starobinsky model as compared to model $R^{\beta}$ and model $\alpha$-Starobinsky. Hence, power law $\alpha$-Starobinsky model is preferred over the other two models.
Generalizations of Starobinsky inflation have both theoretical and phenomenological implications. As the original Starobinsky model is now disfavored at the $2\sigma$ level by the recent ACT observations ACT:2025tim , it highlights the need to study various generalizations of Starobinsky inflation. Moreover, the potential for various generalizations of Starobinsky inflation can be incorporated in the framework of supergravity, which can have several implications for particle physics phenomenology Ellis:2016ipm ; Ellis:2017jcp ; Ellis:2019jha ; Ellis:2019opr .